OpenStudy (anonymous):

Put the following radical expression into simplified form. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\frac{ 4\sqrt{12x^4} }{ \sqrt{3x} }\]

OpenStudy (anonymous):

I don't know if I did it right... I factored \[\sqrt{12x^4}=\sqrt{4}\sqrt{3}=2\sqrt{3}\] Does the x^2 remain in the radical or outside?

OpenStudy (acxbox22):

Outside

OpenStudy (anonymous):

Then I multiplied it by the 4 & got... \[4*2\sqrt{3}=8\sqrt{3}\]

OpenStudy (mathmale):

There are several ways in which you could carry out this simplification. You're off to a good start. To answer your question, that x^2, the square root of x^4, would be written outside.

OpenStudy (mathmale):

so far, so good!

OpenStudy (anonymous):

Do i multiply the top & bottom by \[\sqrt{3x}\] to get rid of the bottom square root?

OpenStudy (mathmale):

Here's a slightly different approach:\[\frac{ 4\sqrt{12x^4} }{ \sqrt{3x} }=\frac{ 4\sqrt{3}\sqrt{4} \sqrt{x^4}}{ \sqrt{3}\sqrt{x}}\]

OpenStudy (mathmale):

Clearly you can cancel that Sqrt(3) in numerator and denominator, and reduce Sqrt(4) to 2. What do you think would be the next step?

OpenStudy (mathmale):

How would you reduce\[\frac{ 4\sqrt{3}\sqrt{4} \sqrt{x^4}}{ \sqrt{3}\sqrt{x}}=\frac{ 8\sqrt{3}x^2 }{ \sqrt{3}x ^{1/2} }\]

OpenStudy (anonymous):

\[\frac{ 8x^2\sqrt{3} }{ 3x }\] Or does the 3 in the denominator remain in a radical?

OpenStudy (mathmale):

Cookie: Cancel the Sqrt(3) that appears in both numerator and denominator:\[\frac{ 8\sqrt{3}x^2 }{ \sqrt{3}x ^{1/2} }\] and then use the rule \[\frac{ x^a }{ x^b }=x ^{a-b}\]

OpenStudy (anonymous):

How do i cancel the top?

OpenStudy (anonymous):

@Miracrown

OpenStudy (mathmale):

\[\frac{ 8\sqrt{3}x^2 }{ \sqrt{3}x ^{1/2} }\] becomes \[\frac{ 8x^2 }{ x ^{1/2} }\]after we've cancelled that Sqrt(3). Then,\[\frac{ x^2 }{ x ^{1/2} }\]

OpenStudy (mathmale):

may be re-written as \[x ^{2-\frac{ 1 }{ 2 }}=?\]

OpenStudy (mathmale):

and so the final answer is\[8x ^{2-\frac{ 1 }{ 2 }}=?\]

OpenStudy (mathmale):

Great working with you! But i need to get off the 'Net now. Good night!

OpenStudy (anonymous):

Goodnight

OpenStudy (anonymous):

8x^3/2